Parth (parthkohli):

A troll suddenly joins OpenStudy and he/she comes to know that a capital i and a lowercase L look the same. How many troll accounts can he/she make with them if the character limit is 10? For example: LiLLiLLLii iLLLLLiiiL LiL Lii LiiiLLL etc. Question 2: He/she comes to know that the 1st letter can't be L because all of them starting with an L are taken. What now?

OpenStudy (aravindg):

pls report abuse for the troll if u see one :P

Parth (parthkohli):

lol no it's a question

ganeshie8 (ganeshie8):

length has to be 10 ? or < 10 is also ok ? he can create \[2^{10}\] unique accounts of 10 character length each b. half of above

Parth (parthkohli):

Yes, he can also create accounts with length of 9,8 and so on..

Parth (parthkohli):

I guess the answer is pretty obvious from now.

Parth (parthkohli):

Do you guys want the solution?

Parth (parthkohli):

Hint: We'll add the combinations for different number of letters ^_^

ganeshie8 (ganeshie8):

\[\sum_{n=1}^{10} 2^{10}\]

Parth (parthkohli):

\(\Large \sum_{k = 1} ^{10} 2^k\) In Summation notation

Parth (parthkohli):

Yeah exactly.

OpenStudy (slaaibak):

2(2^10-1)

Parth (parthkohli):

Yeah, as we know, the answer to the second one would be half of that, so we can say: \(\Huge \sum_{k = 1} ^{10} {2^{k} \over 2} = 2^0 + 2^1 +...2^9\)

ganeshie8 (ganeshie8):

@slaaibak thats much better... looks cute...

Parth (parthkohli):

@AccessDenied when I type the same thing with equation editor, it comes out under the sigma, but when I type it out with hand, it's on the right of the sigma.

Parth (parthkohli):

by hand*

OpenStudy (accessdenied):

Are you using parentheses when you type it out? ( )

OpenStudy (accessdenied):

By that I mean the actual in-line open/close tags

Parth (parthkohli):

\(\Large\sum_{k = 1}\) I'm using this \sum_{k = 1}^{10}

ganeshie8 (ganeshie8):

\[\sum_{n=1}^{10} 2^{10} = 2(2^{10}-1)\]

Parth (parthkohli):

Yeah, inline

Parth (parthkohli):

Oh yeah I got it

Parth (parthkohli):

It doesn't work with the inline tags

OpenStudy (accessdenied):

Yeah, when you use in-line, it does that. You have to use a command to fix that, I think its \limits?

OpenStudy (accessdenied):

Yeah, \sum \limits _{... } ^{... } and so on

Parth (parthkohli):

Where else does the \limits command work?

Parth (parthkohli):

\(\LARGE \sum \limits_{n = 1}^{10}2^n\) I see. Works beautifully :D

Parth (parthkohli):

\(\LARGE\prod \limits_{n = 1}^{10} 2^n \) Works in the product one as well.

OpenStudy (accessdenied):

I believe whenever you have a situation with something being written on top / on bottom of text, it can work. Like, \prod \limits, \lim \limits, and \int \limits

Parth (parthkohli):

I see.

OpenStudy (anonymous):

I love it xD

Parth (parthkohli):

See all lol

Parth (parthkohli):

<--- on the left